A person with myopia has a lens-to-retina distance of 2.0 cm and a far-point of 0.95 m.

What is the minimum optical power of this person's eye? (to 2 s.f and in D)
Answer #1

Here we have given that,

Condition - myopia (nearsightedness)

lens-to-retina distance = 2.0 cm

far-point = 0.95 m.

For the minimum optical power of this person's eye we can use,

Pf = 1/f = 1/v -1/u = 100/2 + 1/0.95 = 51.05263 D

Now for the normal operation of eye Pf' = 1/f'

= 1/0.02 = 50

So that minimum optical power of the person is given as,

Pg = Pf-Pf' = 1.05263 D